For Analytical Chemistry, you need to prepare a buffered mobile phase for an upcoming nicotine HPLC experiment.  You need to prepare a solution buffered to a pH =4.70  and have 1.000 L of a 0.050 M benzoic acid solution.  Since you do not have any benzoate salts available, you decide to make your solution via pH adjustment with a standardized 0.20 M NaOH solution you have left over from a previous experiment.  Calculate the volume of NaOH solution needed to make this adjustment.
Answer #1

pH = pKa + log([PhCOONa]/[PhCOOH]) = 4.70

log([PhCOONa]/[PhCOOH]) = 4.70 - 4.202 = 0.498

[PhCOONa]/[PhCOOH] = 100.498 = 3.1477

Let V ml of 0.20M NaOH is added to 1.000L of 0.050M benzoic acid .

Millimole of benzoic acid taken = 0.050mol/L × 1000ml = 50mmol

Millimole of NaOH taken = 0.20mol/L×V ml = 0.2×V mmol

Ratio of sodium benzoate : benzoic acid = 3.1477

0.2V/(50 - 0.2V) = 3.1477

V = 189.726

Volume of NaOH needs to be added = 189.7 ml